### v3

@@ Abstract
[…]
Research by Da Lozzo, Liotta, and Patrignani was supported, in part, by MUR of Italy (PRIN Project no. 2022ME9Z78  – NextGRAAL and PRIN Project no. 2022TS4Y3N – EXPAND). Research by Liotta was supported in part by MUR PON Proj. ARS01_00540. Research by Rutter was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 541433306. Research by Biedl was supported by NSERC FRN RGPIN-2020-03958.
@@ 1 Introduction
Graph drawing “beyond planarity” studies the combinatorial and algorithmic questions related to representations of graphs where edges can cross but some crossing configurations are forbidden. Depending on the forbidden crossing configuration, different beyond-planar types of drawings can be defined including, for example, RAC, $k$-planar, fan planar, and quasi planar drawings. See [20, 34, 37] for surveys and books.
While most of the literature about beyond planar graph drawing has focused on undirected graphs (one of the few exceptions being [2, 3] which studies RAC upward drawings), we study upward $k$-planar drawings of acyclic digraphs (DAGs), i.e., drawings of DAGs where the edges monotonically increase in $y$-direction and each edge can be crossed at most $k$ times. The minimum $k$ such that a DAG admits an upward $k$-planar drawing is called its upward local crossing number. We focus on values of $k=1,2$ and investigate both combinatorial properties and complexity questions. Our research is motivated by the observation that well-known DAGs that are not upward planar, i.e., not upward 00-planar, do admit a drawing where every edge is crossed at most a constant number of times; see, e.g., Figure 1.
@@
  • A graph is an outerpath if it has a planar drawing in which each vertex is incident to the outer face and the internal faces induce a path in the dual graph. Papakostas [40] observed that there is a directed acyclic 8-vertex outerpath that is not upward-planar (see Figure 3(a)). We strengthen this observation by showing that there exists a directed acyclic fan (that is, a very specific outerpath) that has no upward-planar drawing (Proposition 3.1). On the other hand, we show that every directed acyclic outerpath is upward 2-planar (Theorem 4.6) and that the upward local crossing number is quadratic in the bandwidth (Theorem 4.1). However, the upward local crossing number of bipartite outerplanar DAGs (Theorem 3.3), bipartite DAGs with bounded pathwidth (Corollary 3.7), and cubic DAGs (Proposition 3.8) is in general unbounded.
  • We show that upward $1$-planarity testing is NP-complete, even for graph families where upward planarity testing can be solved in polynomial time. These include: single-source single-sink series-parallel DAGs with a fixed rotation system; single-source two-sink series-parallel DAGs where the rotation system is not fixed; and single-source single-sink DAGs without fixed rotation system that can be obtained from a $K_{4}$ by replacing the edges with series-parallel DAGs (Theorem 5.3).
  • Finally, following a common trend in the study of beyond planar graph representations, we consider the outer model, in which all vertices are required to lie on a common face while maintaining the original requirements [20, 34, 37]. We prove that testing whether a single-source DAG admits an upward outer-1-planar drawing can be done in linear time (Theorem 6.1).
@@
A drawing of a graph is monotone if all edges are drawn monotone with respect to some direction, e.g., a drawing is y-monotone or upward, if each edge intersects each horizontal line at most once. The corresponding crossing number is introduced and studied in [42, 25]. Schaefer [41] mentions the upward crossing number and the local crossing number but not their combination. Schaefer [41, p. 64] also showed that a drawing with the minimum number of crossings per edge can require incident edges that cross. The edges of the provided 4-planar example graph can be oriented such that the resulting directed graph admits an upward 4-planar drawing. Thus, also an upward drawing that achieves the minimum local crossing number can require incident edges that cross. Also, the so-called strong Hanani–Tutte theorem carries over to directed graphs: Fulek et al. [25, Theorem 3.1] showed that every undirected graph that has a monotone drawing where any pair of independent edges crosses an even number of times also has a planar monotone drawing with the same vertex positions. This implies that in any upward drawing of a graph that is not upward-planar there must be a pair of independent edges that crosses an odd number of times.
Upward drawings of directed acyclic graphs have been studied in the context of (upward) book embeddings. In that model the vertices are drawn on a vertical line (a spine) following a topological order of the graph, while all edges are pointing upwards. To reduce the edge crossings, edges are partitioned into the fewest number of crossing-free subsets (pages). Studying upward book embeddings is a popular topic, which is usually centered around determining the smallest number of pages for various graph classes [31, 23, 24, 39, 36] or deciding whether a graph admits an upward drawing with a given number of pages [11, 9, 10, 5, 6]. Our model is related to topological book embeddings [38, 29], which are a relaxed version of 2-page book embeddings in which edges are allowed to cross the spine. While papers about topological book embeddings insist on planar drawings and minimize the number of spine crossings, we do allow crossings and want to bound the maximum number of crossings per edge (ignoring the spine).
@@ 3 Lower Bounds
We start with a negative result that shows that even very special directed acyclic outerpaths may not admit upward-planar drawings, thus strengthening Papakostas’ observation [40].
Proposition 3.1.
@@
Proof 3.2.
Consider the 7-vertex fan $F$ depicted in Figure 2(a). Suppose for a contradiction that $F$ is upward planar, that is, $F$ admits an upward planar drawing $\Gamma$. Let $c$ be the central vertex of $F$. We assume that $c$ is placed at the origin. We say that a triangle of $F$ is positive (negative, respectively) if the corresponding region of the plane in $\Gamma$ contains the point $(\varepsilon,0)$ ($(-\varepsilon,0)$, respectively) for a sufficiently small value $\varepsilon>0$. The triangles that have one vertex below $c$ and one vertex above $c$ (namely $t_{1}=\triangle cv_{1}v_{2}$, $t_{3}=\triangle cv_{3}v_{4}$, and $t_{5}=\triangle cv_{5}v_{6}$) are either positive or negative.
@@
If both $t_{1}$ and $t_{5}$ are positive, then one must contain the other in $\Gamma$, say, $t_{1}$ contains $t_{5}$; see Figure 2(b). But then vertices $v_{3}$ and $v_{4}$ must also lie inside $t_{1}$. If both lie inside $t_{5}$, then the edge $(v_{3},v_{2})$ intersects an edge of $t_{5}$. If one of them lies inside $t_{5}$ and one does not, then the edge $(v_{4},v_{3})$ intersects an edge of $t_{5}$. So both must lie outside $t_{5}$. But $v_{4}$ lies on one hand above $v_{5}$ and on the other hand below $c$ and, thus, below $v_{6}$. So the edge $(v_{4},c)$ intersects the edge $(v_{5},v_{6})$. (If $t_{5}$ contains $t_{1}$, the edge $(c,v_{3})$ intersects the edge $(v_{1},v_{2})$.)
By symmetry, not both $t_{1}$ and $t_{5}$ can be negative, so exactly one of $t_{1}$ and $t_{5}$ must be negative, say, $t_{1}$; see Figure 2(c). Now first assume that $t_{3}$ is positive. Due to edge $(v_{3},v_{2})$, vertex $v_{3}$ must be outside $t_{5}$, so $t_{3}$ cannot be inside $t_{5}$. On the other hand, $t_{3}$ cannot contain $t_{5}$ because $v_{4}$ is above $v_{5}$. Hence $t_{3}$ intersects $t_{5}$. Finally, assume that $t_{3}$ is negative. Due to edge $(v_{5},v_{4})$, vertex $v_{4}$ must be outside $t_{1}$, so $t_{3}$ cannot be inside $t_{1}$. On the other hand, $t_{3}$ cannot contain $t_{1}$ because $v_{3}$ is below $v_{2}$. Hence $t_{3}$ intersects $t_{1}$.
@@
Proof 3.4.
The bipartite graph $G_{0}$ in Figure 3(a) is not upward planar [40]. For $\ell\geq 1$, we construct $G_{\ell}$ from $G_{\ell-1}$ by adding a $3$-edge path on every outer edge of the graph. Figure 3(b) shows $G_{2}$. The maximum degree of $G_{\ell}$ is $\Delta_{\ell}=2\ell+3$. The number of vertices is $n_{\ell}=8+\sum_{i=1}^{\ell}8\cdot 3^{i-1}\cdot 2=8\cdot 3^{\ell}$.
@@
We now determine a cycle $C$ of $G_{\ell}$ that is crossed by $e$ an odd number of times and does also not contain any end vertex of $e$. If $e^{\prime}$ is an inner edge, then we take the outer path $P$ of $G_{0}$ that connects the ends of $e^{\prime}$ and does not contain $e$; this is not crossed by $e$ due to our choice of $e^{\prime}$. Moreover, since $e^{\prime}$ and $e$ are independent, it follows that $P$ and $e$ do not share a vertex. Let $C$ be the concatenation of $P$ and $e^{\prime}$. In this case $C$ has length at most six. See Figure 4(a).
If $e^{\prime}$ is an outer edge of $G_{0}$, we do the following: We start with the path $P$ of length three that was added for $e^{\prime}$. Since $e$ is an edge of $G_{0}$ and $e$ and $e^{\prime}$ are independent, it follows that $P$ and $e$ do not share a vertex. If $P$ contains an edge that is crossed an odd number of times by $e$ then we continue with such an edge instead of $e^{\prime}$. More precisely, let $e_{1}=e^{\prime}$ and initialize $i=1$. Let $P_{i}$ be the path of length three that was added for $e_{i}$. While $P_{i}$ contains an edge that is crossed an odd number of times by $e$, let $e_{i+1}$ be such an edge, let $P_{i+1}$ be the path of length three that was added for $e_{i}$, and increase $i$ by one. Since $e$ is crossed at most $k$ times, this process stops at some $i\leq k$.Let $C$ be the cycle that is composed of $P_{i}$ and $e_{i}$. See Figure 4(b). In this case $C$ has length four. Moreover $C$ shares at most the end vertices of $e^{\prime}$ with $G_{0}$. Thus, since $e$ is an edge of $G_{0}$ and $e$ and $e^{\prime}$ are independent, it follows that $C$ and $e$ do not share a vertex.
Cycle $C$ might cross itself. However, it divides the plane into cells. Since $e$ crosses $C$ an odd number of times, it follows that the end vertices of $e$ must be in different cells of the plane. This means that not only $e$ but also the $\ell$ edge-disjoint paths that were added on top of $e$ have to cross $C$. Observe that none of these paths contains a vertex of $C$. But $C$ contains at most six edges, each of which can be crossed at most $k$ times. This is impossible if $\ell\geq 6k$. Hence, if there is an upward $k$-planar drawing then $\ell<6k$, which means that $k>\ell/6$.
We now show that if we expand the graph class beyond outerplanar graphs, then we get a lower bound on the upward local crossing number that is even linear in the number of vertices. The graphs in our construction have pathwidth 2, as opposed to the graphs in Theorem 3.3 whose pathwidth is logarithmic. Observe that a caterpillar, i.e., a tree that can be reduced to a path by removing all degree-1 vertices, has pathwidth 1, and that the pathwidth can increase by at most 1 if we add a vertex with some incident edges or subdivide some edges.
Theorem 3.5.
@@
Proof 3.6.
Let $G_{k}$ be the graph consisting of the four vertices $a$, $b_{1}$, $b_{2}$, and $c$ and the following set of edges and degree-2 vertices (see also Figure 5):
@@
Curve $C_{b}$ uses six edges, therefore among the $16k+$ 1 sources below $a$, there exists one, call it $s$, for which edge $(s,c)$ crosses no edge of $C_{b}$. Since $y(s)<y(a)$, vertex $s$ is outside $C_{b}$, and so the entire edge $(s,c)$ is outside $C_{b}$, except at the endpoint $c$. In particular, among the three edges $(d_{1},c)$, $(d_{2},c)$, and $(s,c)$ that are incoming at $c$, edge $(s,c)$ is either leftmost or rightmost (but cannot be the middle one). We assume here that $(s,c)$ is rightmost, the other case is symmetric. Write $\{p,q\}=\{1,2\}$ such that the left-to-right order of incoming edges at $c$ is $(d_{p},c)$, $(d_{q},c)$, $(s,c)$. In Figure 5, we have $p=1$ and $q=2$.
Edge $(s,a)$ is also outside $C_{b}$, except perhaps at endpoint $a$, since it uses smaller $y$-coordinates. Let $s^{\prime}$ be the topmost intersection point of $(s,a)$ and $(s,c)$. Then there are no crossings in the curve $C_{a}$ formed by the directed paths $\langle s^{\prime},a,b_{p},d_{p},c\rangle$ and $\langle s^{\prime},c\rangle$. By our choice of $p$ and $q$, vertex $d_{q}$ is inside $C_{a}$, and so is the entire path $\langle a^{\prime},b_{q},d_{q},c\rangle$, except at the ends since it is part of $C_{b}$. In particular, $b_{q}$ is inside $C_{a}$, whereas, for $j\in[4k+1]$, $t_{q}^{(j)}$ is outside $C_{a}$ due to $y(c)<y(t_{q}^{(j)})$. It follows that one of the four edges $(a,b_{p})$, $(b_{p},d_{p})$, $(d_{p},c)$ and $(s,c)$ must be crossed at least $k+1$ times by edges from $b_{q}$ to the sinks above it. Thus, the drawing was not $k$-planar, a contradiction.
The graphs that we constructed in the proof of Theorem 3.5 are not bipartite, but one can make them bipartite by subdividing all edges once. This at best cuts the local crossing number in half, increases the pathwidth by at most 1, and yields the following result.
Corollary 3.7.
@@
Proof 3.9.
The crossing number of a random cubic graph with $n$ vertices is expected to be at least $cn^{2}$ for some absolute constant $c>0$ [21], and thus there exist graphs yielding this bound. By the pigeon-hole principle, such a graph contains an edge with $\Omega(n)$ crossings among its $\Theta(n)$ edges. Impose arbitrary acyclic edge directions.

### v4

@@ 1 Introduction
Graph drawing “beyond planarity” studies the combinatorial and algorithmic questions related to representations of graphs where edges can cross but some crossing configurations are forbidden. Depending on the forbidden crossing configuration, different beyond-planar types of drawings can be defined including, for example, RAC, $k$-planar, fan planar, and quasi planar drawings (see Didimo, Liotta, and Montecchiani, 2019; Hong and Tokuyama, 2020; Kobourov, Liotta, and Montecchiani, 2017, for surveys and books).
While most of the literature about beyond planar graph drawing has focused on undirected graphs (one of the few exceptions being the paper of Angelini, Cittadini, Didimo, Frati, Di Battista, Kaufmann, and Symvonis (2011), which studies RAC upward drawings), we study upward $k$-planar drawings of acyclic digraphs (DAGs), i.e., drawings of DAGs where the edges monotonically increase in $y$-direction and each edge can be crossed at most $k$ times. The minimum $k$ such that a DAG admits an upward $k$-planar drawing is called its upward local crossing number. We focus on values of $k=1,2$ and investigate both combinatorial properties and complexity questions. Our research is motivated by the observation that well-known DAGs that are not upward planar, i.e., not upward 00-planar, do admit a drawing where every edge is crossed at most a constant number of times; see, e.g., Fig. 1.
@@
  • A graph is an outerpath if it has a planar drawing in which each vertex is incident to the outer face and the internal faces induce a path in the dual graph. Papakostas (1994) observed that there is a directed acyclic 8-vertex outerpath that is not upward-planar (see Fig. 3(a)). We strengthen this observation by showing that there exists a directed acyclic fan (that is, a very specific outerpath) that has no upward-planar drawing (Proposition 1). On the other hand, we show that every directed acyclic outerpath is upward 2-planar (Theorem 4) and that the upward local crossing number is quadratic in the bandwidth (Theorem 3). However, the upward local crossing number of bipartite outerplanar DAGs (Theorem 1), bipartite DAGs with bounded pathwidth (Corollary 1), and cubic DAGs (Proposition 2) is in general unbounded.
  • We show that upward $1$-planarity testing is NP-complete, even for graph families where upward planarity testing can be solved in polynomial time. These include: single-source single-sink series-parallel DAGs with a fixed rotation system; single-source two-sink series-parallel DAGs where the rotation system is not fixed; and single-source single-sink DAGs without fixed rotation system that can be obtained from a $K_{4}$ by replacing the edges with series-parallel DAGs (Theorem 6).
  • Finally, following a common trend in the study of beyond planar graph representations, we consider the outer model, in which all vertices are required to lie on a common face while maintaining the original requirements (Didimo, Liotta, and Montecchiani, 2019; Hong and Tokuyama, 2020; Kobourov, Liotta, and Montecchiani, 2017). We prove that testing whether a single-source DAG admits an upward outer-1-planar drawing can be done in linear time (Theorem 7).
@@
A drawing of a graph is monotone if all edges are drawn monotone with respect to some direction, e.g., a drawing is y-monotone or upward, if each edge intersects each horizontal line at most once. The corresponding crossing number is introduced and studied by Valtr (2005) and Fulek, Pelsmajer, Schaefer, and Štefankovič (2013). Schaefer (2024) mentions the upward crossing number and the local crossing number but not their combination. Schaefer (2024, p. 64) also showed that a drawing with the minimum number of crossings per edge can require incident edges that cross. The edges of the provided 4-planar example graph can be oriented such that the resulting directed graph admits an upward 4-planar drawing. Thus, also an upward drawing that achieves the minimum local crossing number can require incident edges that cross. Also, the so-called strong Hanani–Tutte theorem carries over to directed graphs: Fulek et al. (2013, Theorem 3.1) showed that every undirected graph that has a monotone drawing where any pair of independent edges crosses an even number of times also has a planar monotone drawing with the same vertex positions. This implies that in any upward drawing of a graph that is not upward-planar there must be a pair of independent edges that crosses an odd number of times.
Upward drawings of directed acyclic graphs have been studied in the context of (upward) book embeddings. In that model the vertices are drawn on a vertical line (a spine) following a topological order of the graph, while all edges are pointing upwards. To reduce the edge crossings, edges are partitioned into the fewest number of crossing-free subsets (pages). Studying upward book embeddings is a popular topic, which is usually centered around determining the smallest number of pages for various graph classes (Heath, Pemmaraju, and Trenk, 1999; Frati, Fulek, and Ruiz-Vargas, 2011, 2013; Nöllenburg and Pupyrev, 2023; Jungeblut, Merker, and Ueckerdt, 2023) or deciding whether a graph admits an upward drawing with a given number of pages (Binucci, Da Lozzo, Di Giacomo, Didimo, Mchedlidze, and Patrignani, 2023; Bhore, Da Lozzo, Montecchiani, and Nöllenburg, 2023; Bekos, Da Lozzo, Frati, Gronemann, Mchedlidze, and Raftopoulou, 2023). Our model is related to topological book embeddings (Mchedlidze and Symvonis, 2009; Giordano, Liotta, Mchedlidze, Symvonis, and Whitesides, 2015), which are a relaxed version of 2-page book embeddings in which edges are allowed to cross the spine. While papers about topological book embeddings insist on planar drawings and minimize the number of spine crossings, we do allow crossings and want to bound the maximum number of crossings per edge (ignoring the spine).
@@ 3 Lower Bounds
We start with a negative result that shows that even very special directed acyclic outerpaths may not admit upward-planar drawings, thus strengthening Papakostas’ observation (1994).
Proposition 1.
@@
Proof.
Consider the 7-vertex fan $F$ depicted in Fig. 2(a). Suppose for a contradiction that $F$ is upward planar, that is, $F$ admits an upward planar drawing $\Gamma$. Let $c$ be the central vertex of $F$. We assume that $c$ is placed at the origin. We say that a triangle of $F$ is positive (negative, respectively) if the corresponding region of the plane in $\Gamma$ contains the point $(\varepsilon,0)$ ($(-\varepsilon,0)$, respectively) for a sufficiently small value $\varepsilon>0$. The triangles that have one vertex below $c$ and one vertex above $c$ (namely $t_{1}=\triangle cv_{1}v_{2}$, $t_{3}=\triangle cv_{3}v_{4}$, and $t_{5}=\triangle cv_{5}v_{6}$) are either positive or negative.
@@
If both $t_{1}$ and $t_{5}$ are positive, then one must contain the other in $\Gamma$, say, $t_{1}$ contains $t_{5}$; see Fig. 2(b). But then vertices $v_{3}$ and $v_{4}$ must also lie inside $t_{1}$. If both lie inside $t_{5}$, then the edge $(v_{3},v_{2})$ intersects an edge of $t_{5}$. If one of them lies inside $t_{5}$ and one does not, then the edge $(v_{4},v_{3})$ intersects an edge of $t_{5}$. So both must lie outside $t_{5}$. But $v_{4}$ lies on one hand above $v_{5}$ and on the other hand below $c$ and, thus, below $v_{6}$. So the edge $(v_{4},c)$ intersects the edge $(v_{5},v_{6})$. (If $t_{5}$ contains $t_{1}$, the edge $(c,v_{3})$ intersects the edge $(v_{1},v_{2})$.)
By symmetry, not both $t_{1}$ and $t_{5}$ can be negative, so exactly one of $t_{1}$ and $t_{5}$ must be negative, say, $t_{1}$; see Fig. 2(c). Now first assume that $t_{3}$ is positive. Due to edge $(v_{3},v_{2})$, vertex $v_{3}$ must be outside $t_{5}$, so $t_{3}$ cannot be inside $t_{5}$. On the other hand, $t_{3}$ cannot contain $t_{5}$ because $v_{4}$ is above $v_{5}$. Hence $t_{3}$ intersects $t_{5}$. Finally, assume that $t_{3}$ is negative. Due to edge $(v_{5},v_{4})$, vertex $v_{4}$ must be outside $t_{1}$, so $t_{3}$ cannot be inside $t_{1}$. On the other hand, $t_{3}$ cannot contain $t_{1}$ because $v_{3}$ is below $v_{2}$. Hence $t_{3}$ intersects $t_{1}$. ∎
@@
Proof.
The bipartite graph $G_{0}$ in Fig. 3(a) is not upward planar (Papakostas, 1994). For $\ell\geq 1$, we construct $G_{\ell}$ from $G_{\ell-1}$ by adding a $3$-edge path on every outer edge of the graph. Fig. 3(b) shows $G_{2}$. The maximum degree of $G_{\ell}$ is $\Delta_{\ell}=2\ell+3$. The number of vertices is $n_{\ell}=8+\sum_{i=1}^{\ell}8\cdot 3^{i-1}\cdot 2=8\cdot 3^{\ell}$.
@@
We now determine a cycle $C$ of $G_{\ell}$ that is crossed by $e$ an odd number of times and does also not contain any end vertex of $e$. If $e^{\prime}$ is an inner edge, then we take the outer path $P$ of $G_{0}$ that connects the ends of $e^{\prime}$ and does not contain $e$; this is not crossed by $e$ due to our choice of $e^{\prime}$. Moreover, since $e^{\prime}$ and $e$ are independent, it follows that $P$ and $e$ do not share a vertex. Let $C$ be the concatenation of $P$ and $e^{\prime}$. Note that the length of $C$ is at most six; see Fig. 4(a).
If $e^{\prime}$ is an outer edge of $G_{0}$, we do the following: We start with the path $P$ of length three that was added for $e^{\prime}$. Since $e$ is an edge of $G_{0}$ and $e$ and $e^{\prime}$ are independent, it follows that $P$ and $e$ do not share a vertex. If $P$ contains an edge that is crossed an odd number of times by $e$ then we continue with such an edge instead of $e^{\prime}$. More precisely, let $e_{1}=e^{\prime}$ and initialize $i=1$. Let $P_{i}$ be the path of length three that was added for $e_{i}$. While $P_{i}$ contains an edge that is crossed an odd number of times by $e$, let $e_{i+1}$ be such an edge, let $P_{i+1}$ be the path of length three that was added for $e_{i}$, and increase $i$ by one. Since $e$ is crossed at most $k$ times, this process stops at some $i\leq k$. Let $C$ be the cycle that is composed of $P_{i}$ and $e_{i}$. See Fig. 4(b). In this case $C$ has length four. Moreover $C$ shares at most the end vertices of $e^{\prime}$ with $G_{0}$. Thus, since $e$ is an edge of $G_{0}$ and $e$ and $e^{\prime}$ are independent, it follows that $C$ and $e$ do not share a vertex.
Cycle $C$ might cross itself. However, it divides the plane into cells. Since $e$ crosses $C$ an odd number of times, it follows that the end vertices of $e$ must be in different cells of the plane. This means that not only $e$ but also the $\ell$ edge-disjoint paths that were added on top of $e$ have to cross $C$. Observe that none of these paths contains a vertex of $C$. But $C$ contains at most six edges, each of which can be crossed at most $k$ times. This is impossible if $\ell\geq 6k$. Hence, if there is an upward $k$-planar drawing then $\ell<6k$, which means that $k>\ell/6$. ∎
We now show that if we expand the graph class beyond outerplanar graphs, then we get a lower bound on the upward local crossing number that is even linear in the number of vertices. The graphs in our construction have pathwidth 2, as opposed to the graphs in Theorem 1 whose pathwidth is logarithmic. Observe that a caterpillar, i.e., a tree that can be reduced to a path by removing all degree-1 vertices, has pathwidth 1, and that the pathwidth can increase by at most 1 if we add a vertex with some incident edges or subdivide some edges.
Theorem 2.
@@
Proof.
Let $G_{k}$ be the graph consisting of the four vertices $a$, $b_{1}$, $b_{2}$, and $c$ and the following set of edges and degree-2 vertices (see also Fig. 5):
@@
Curve $C_{b}$ uses six edges, therefore among the $16k+$ 1 sources below $a$, there exists one, call it $s$, for which edge $(s,c)$ crosses no edge of $C_{b}$. Since $y(s)<y(a)$, vertex $s$ is outside $C_{b}$, and so the entire edge $(s,c)$ is outside $C_{b}$, except at the endpoint $c$. In particular, among the three edges $(d_{1},c)$, $(d_{2},c)$, and $(s,c)$ that are incoming at $c$, edge $(s,c)$ is either leftmost or rightmost (but cannot be the middle one). We assume here that $(s,c)$ is rightmost, the other case is symmetric. Write $\{p,q\}=\{1,2\}$ such that the left-to-right order of incoming edges at $c$ is $(d_{p},c)$, $(d_{q},c)$, $(s,c)$. In Fig. 5, we have $p=1$ and $q=2$.
Edge $(s,a)$ is also outside $C_{b}$, except perhaps at endpoint $a$, since it uses smaller $y$-coordinates. Let $s^{\prime}$ be the topmost intersection point of $(s,a)$ and $(s,c)$. Then there are no crossings in the curve $C_{a}$ formed by the directed paths $\langle s^{\prime},a,b_{p},d_{p},c\rangle$ and $\langle s^{\prime},c\rangle$. By our choice of $p$ and $q$, vertex $d_{q}$ is inside $C_{a}$, and so is the entire path $\langle a^{\prime},b_{q},d_{q},c\rangle$, except at the ends since it is part of $C_{b}$. In particular, $b_{q}$ is inside $C_{a}$, whereas, for $j\in[4k+1]$, $t_{q}^{(j)}$ is outside $C_{a}$ due to $y(c)<y(t_{q}^{(j)})$. It follows that one of the four edges $(a,b_{p})$, $(b_{p},d_{p})$, $(d_{p},c)$ and $(s,c)$ must be crossed at least $k+1$ times by edges from $b_{q}$ to the sinks above it. Thus, the drawing was not $k$-planar, a contradiction. ∎
The graphs that we constructed in the proof of Theorem 2 are not bipartite, but one can make them bipartite by subdividing all edges once. This at best cuts the local crossing number in half, increases the pathwidth by at most 1, and yields the following result.
Corollary 1.
@@
Proof.
The crossing number of a random cubic graph with $n$ vertices is expected to be at least $cn^{2}$ for some absolute constant $c>0$ (Dujmović, Kawarabayashi, Mohar, and Wood, 2008), and thus there exist graphs yielding this bound. By the pigeon-hole principle, such a graph contains an edge with $\Omega(n)$ crossings among its $\Theta(n)$ edges. Impose arbitrary acyclic edge directions. ∎